OpenStudy (anonymous):

i solved (2x+4)/(x-2) subtracted from (x+3)/(x-2) and got x+7/x-2. am i right?

OpenStudy (experimentx):

i guess you are wrong, check yourself from now on http://www.wolframalpha.com/input/?i=%28x%2B3%29%2F%28x-2%29+-+%282x%2B4%29%2F%28x-2%29

OpenStudy (anonymous):

all those options aren't even possible answers as my options.

OpenStudy (experimentx):

(x+3)/(x-2) - (2x+4)/(x-20) =((x+3)-(2x+4))/(x-2) =(x-2x +3 -4)/(x-2) (-x-1)/(x-2)

OpenStudy (anonymous):

thats not one of my options. i think you mean x+1/x-2